# Public Infrastructure Investments for Space Market Development

### v1

@@ Abstract
Advanced space technology systems often face high fixed costs, can serve limited non-government demand, and are significantly driven by non-market motivations. While increased entrepreneurial activity and national ambitions in space have encouraged planners at public space agencies to develop markets around such systems, the very factors that make the recent growth of the space economy so remarkable also challenge planners’ efforts to develop and sustain markets for space-related goods and services. I propose a graphical framework to visualize the number of competitors a market can sustain as a function of the industry’s cost structure; the distribution of government support across direct purchases, direct investments, and shared infrastructure; and the magnitude of non-government demand. Building on public goods theory, the framework shows how marginal dollars invested in shared infrastructure can create non-rival benefits supporting more competitors per dollar than direct purchases or subsidies. I demonstrate the framework with a stylized application inspired by NASA’s Commercial LEO Destinations program. Under cost and demand conditions consistent with public data, independent stations generate industry-wide losses of \$355 million annually, while shared core infrastructure enables industry-wide profits of \$154 million annually. I also outline key directions for future research on public investment and market development strategies for advanced technologies.
@@ 1 Introduction
Public space agencies are increasingly interested in developing sustainable markets for advanced space technologies. Several have adopted market development goals, in some cases seeking to transition from direct provision to commercial procurement of space services (NASA, 2019; European Commission, 2025; ESA, 2025a; Cabinet Office, Government of Japan, 2024; EY Global, 2024) Yet a basic economic question—how should a public space agency invest in space activities to develop and sustain markets for space-related goods and services—remains underexplored. While systems like the International Space Station (ISS) demonstrate clear technical capabilities—from scientific research across domains as diverse as biology (Castro-Wallace et al., 2017), astronomy (Magalhães et al., 2021), and economics (Henderson et al., 2012) to emerging commercial applications like ZBLAN glass production (Werner, 2024)—the transition from government-funded capabilities to competitive markets remains challenging. High fixed costs, long development timelines, and uncertain market demand create conditions where programs to develop and sustain space markets must contend with the risk that the supply side collapses, leaving a monopoly or no market at all. Space technology systems also serve non-market motivations like signaling national achievement or capability (MacDonald, 2017). These goals complicate market development efforts, as market forces alone may not demand sufficiently capable systems.
Space agencies also often have dual objectives: they must procure mission inputs—transportation services, spacecraft components, advanced research facilities—cost-effectively, while also ensuring the continued existence of industrial capabilities that serve limited non-government demand and have high barriers to entry. As space ambitions rise globally while public budgets face greater strain, understanding and clearly communicating principles of pro-competitive public investment in advanced technology systems is an increasingly important task for economists and public space agency officials. In this essay I propose a graphical framework to assist space agency planners in designing advanced technology programs to encourage competition and market development.
The framework enables space agency planners to assess how different distributions of program budgets across direct purchases, direct transfers, and shared infrastructure affect the number of competitors a market can sustain. It supports analysis of what I term “economic architecture.” Just as systems engineering uses “architecture” to describe how technical components fit and work together, a program’s economic architecture determines which and how many technology systems are economically viable. Building on the standard theory of public goods (Samuelson, 1954), the key insight is that marginal dollars invested in shared infrastructure can create non-rival benefits supporting more competitors per dollar than direct purchases or subsidies. By plotting the number of competitors sustainable at different levels of total direct government revenues and industry total costs, the diagram reveals regions where different numbers of firms can earn positive profits, showing precisely when shared infrastructure investments enable competition that direct support mechanisms cannot sustain.
@@
I demonstrate the framework using a stylized application inspired by NASA’s Commercial LEO Destinations program, comparing independent free-flyer versus shared core module architectures for sustaining at least two competing crewed space station operators. Under cost and demand conditions consistent with public data, independent stations generate industry-wide losses of \$355 million annually, while shared core infrastructure enables each firm to earn \$77 million in annual profits. Scenario analysis reveals how external factors like market demand and interest rates can overwhelm program design choices, underscoring market development programs’ vulnerability to conditions beyond agency control.
This work addresses gaps across three research strands. The public economics literature has explored optimal public-private contracts (Laffont and Tirole, 1993; Danau and Vinella, 2015) and developed typologies for space sector partnerships (Kim, 2023), but provides few actionable frameworks for sustaining advanced technology industrial sectors. [^1] Unlike statistical or financial agencies, public space agencies do not usually have internal economic analysis teams that utilize the types of analytical tools developed in the economics literature. NASA’s Chief Economist role was, to my knowledge, one of the few such functions in a public space agency globally (SpaceRef, 2019). This paper is informed by my work with the Chief Economist team within NASA’s Office of Technology, Policy, and Strategy. Among other things, the team developed estimates of non-NASA demand in markets of strategic interest to the agency and assessed economic architectures for existing and proposed programs. All inputs for that demand estimation were derived from publicly available sources. The Chief Economist function was eliminated with the Office of Technology, Policy, and Strategy in early 2025 (Smith, 2025). Techno-economic analyses focus on single system architectures (Proctor et al., 2021; Colvin et al., 2020a; Metzger, 2023) while remaining silent on competitive dynamics or optimal support mechanisms. “Strategic” demand estimation capabilities for nascent space markets remain underdeveloped, with few such assessments existing beyond those produced by think tanks for specific government sponsors (Crane et al., 2017; Colvin et al., 2020b; Triezenberg et al., 2020; MITRE, 2021; Triezenberg et al., 2024). “Strategic” demand estimates need not be econometrically identified, but must be constructible for nascent or non-existent markets within weeks or months, quantify demand across years or decades to within an order of magnitude, and reflect relevant technical system characteristics and their likely evolutions. By integrating insights across these areas into a simple graphical tool, this framework provides planners at space agencies with a systematic approach to assess program design trade-offs.
The structure of this essay is as follows. In section 2 I describe the general framework for assessing economic architectures of market development programs. This framework is generic, though it is best suited to advanced technology systems with high fixed costs, long development timelines, and highly uncertain market demand. In section 3 I apply the framework to private space stations in low-Earth orbit. In section 4 I identify key directions for future research on public investment and market development strategies for advanced technologies, including improved economic measurement, strategic demand estimation, and better understanding of how private investors perceive different revenue sources. I conclude in section 5.
2 A framework for assessing economic architectures of market development programs
@@
Consider a public space agency with annual budget $B$ designing a program to develop sustainable markets for a class of advanced space technology systems. These systems—such as crewed space stations, lunar surface power systems, or orbital debris removal spacecraft—typically exhibit high fixed costs, long development timelines, and uncertain non-government demand, creating conditions where uncoordinated private investment may result in too few competitors or none at all. [^2] As Weinzierl (2018) notes, complementarities between space technologies mean space businesses across the technology stack play stag hunts with each other. This framework assumes the public space agency is a planner capable of structuring technological complementarities between private actors to encourage desired outcomes. I refer to the $N$ firms that develop and utilize the technology systems of interest as “the industry.” I express all economic variables as equivalent annuity units to facilitate comparison across different payment timing structures, assume firms are symmetric in costs and capabilities, and ignore inflation.
@@
Public investment into shared infrastructure produces a flow of shared benefits $Y^{G}$ for the industry according to the production and valuation function $f(G^{S})$. This function measures the extent to which shared infrastructure reduces firms’ costs relative to developing equivalent capabilities independently. For modular technology systems like space stations, shared infrastructure benefits may be approximately linear up to the cost of shareable essential elements—such as life support systems or power generation—then exhibit sharply diminishing returns thereafter. For network systems requiring interoperability, benefits may show increasing returns to scale as more participants connect to the shared network component. The specific functional form depends on the technology systems and business models involved, and could vary across firms. Here, I assume firms uniformly value shared infrastructure at cost: $Y^{G}=G^{S}$.
Firms in the industry earn total revenue $R_{i}$, with $R_{i}^{G}$ earned from the program and $R^{M}_{i}$ earned from “the market.” “The market” captures all sources of demand which are exogenous to the agency—e.g., individuals, households, firms, other public space agencies, and any other public or private entities that may purchase the industry’s output. In specific applications, revenues earned from the agency or the market can be divided between firms according to market share functions $\sigma^{G}_{i}(R^{G})\equiv R^{G}_{i}$ and $\sigma^{M}_{i}(R^{M})\equiv R^{M}_{i}$. Under symmetry each firm receives an equal share of both government and market revenues, so these simplify to $\sigma^{G}_{i}(R^{G})=R^{G}_{i}=R^{G}/N$ and $\sigma^{M}_{i}(R^{M})=R^{M}_{i}=R^{M}/N$. Similarly, direct transfers are simplified to $G_{i}=G^{D}/N$. [^4] Allowing for asymmetry introduces additional modeling choices such as the order of entry or exit, quickly exploding the number of possible cases. Institutional knowledge can guide analysts toward particular asymmetric scenarios.
Total industry revenue is $R=\sum_{i}(R_{i}^{G}+R_{i}^{M})$. Firms in the industry incur gross total cost $X_{i}$ and net total cost $C_{i}=X_{i}-Y^{G}-G_{i}$, and industry’s total cost is $C=\sum_{i}C_{i}$. The gross costs $X_{i}$ represent the costs to develop, deploy, and operate the target advanced technology system. Let $X=\sum_{i}X_{i}$ be the industry-level gross total cost. Under symmetry, $X_{i}=X/N$. While these costs must be controlled, it is widely understood by space technology program managers that there is often a positive relationship between a system’s performance and its cost. For example, Section 4.2.1 of NASA’s Cost Estimating Handbook states: “The objective of a cost-performance trade study is not necessarily to minimize the cost of the system, but to achieve an optimal balance of performance and cost” (NASA Executive Cost Analysis Steering Group, 2015). With suitable caveats, then, an advanced technology system’s cost may be used a proxy for its performance during program design.
@@
Agencies intending to promote industry may allocate some of their budget towards increasing industry profits. A marginal dollar provides greater support to the industry if invested in shared infrastructure than if used for direct purchasing when:
@@
If the public space agency wants to maintain at least $N$ competitors, then it must ensure $R^{G}$, $G^{D}$, and $Y^{G}$—its support for the industry—are sufficient to ensure all $N$ firms earn non-negative profits, conditional on the anticipated level of market demand, $R^{M}$. Formally:
@@
This constraint captures the fundamental trade-off facing market development programs: supporting more competitors requires either reducing individual firm costs (through shared infrastructure investments) or increasing industry revenues (through direct purchases, transfers, or market growth). Programs that violate this constraint will see firms exit, potentially ending in monopoly or no suppliers at all. The agency can use firms’ profitability constraints and its own budget constraint to assess high-level program design trade-offs, particularly how to allocate budget across direct purchases, direct transfers, and shared infrastructure. Questions regarding the timing and conditioning of specific payments or infrastructure access can be investigated by appropriately modeling and discounting the flows of benefits received by the industry when calculating $R^{G}$, $G^{D}$, and $Y^{G}$.
The full architectural possibility set can be visualized by taking intersections of industry total costs and government direct purchases where each firm earns non-negative profits at a given industry size, and then overlaying these regions across different industry sizes. For fixed industry size $N$, these regions are formally described by:
$\{(R^{G},C):~\forall i,~\Pi_{i}(N)\geq 0\},$ (5)
@@
i.e., intersections of superlevel sets for firms’ profitability constraints under the given assumptions. [^5] Similar diagrams are used in managing engineering costs and schedules in advanced technology programs, e.g., Joint Confidence Level plots (NASA Executive Cost Analysis Steering Group, 2015). The resulting diagram can be used to assess how many competitors can be sustained by different economic architectures under consistent assumptions. Direct government purchases from the industry, $R^{G}$, is the portion of agency support that appears on firms’ income statements. Industry total costs, $C=X-Y^{G}-G^{D}$, shows the cost burden firms face after accounting for shared infrastructure benefits and direct transfers. Regions where profits are non-negative for a fixed industry size show combinations of total costs and direct purchases that can sustain that many firms. Within these regions, increases in direct purchases can be traded off against reductions in net industry costs to maintain the same expected level of competition.
To make this concrete, suppose a public space agency is considering two alternative economic architectures for a program intended to sustain at least two firms developing and utilizing certain technology systems. In architecture A the agency will directly transfer all funds to the industry as non-dilutive grants, while in architecture B the agency will evenly divide its budget between direct purchases and shared infrastructure. Industry’s expected market revenues are valued at \$1B/year (i.e., \$0.5B/year/firm), industry gross total costs are \$2B/year with two firms in the industry (i.e., \$1B/year/firm), and the program has \$1B/year to allocate. Suppose also that firms uniformly value the benefits provided by shared infrastructure at cost. [^6] In this example, analysis of condition (3) shows that this assumption implies shared infrastructure better supports the industry than direct purchasing or direct investment at all industry sizes. Table 1 lists the economic parameters for an individual firm under both economic architecture assuming two competitors, with parameter choices selected to illustrate the framework’s key insights about infrastructure versus direct support trade-offs.
I assumed the program intends to sustain at least two firms. As Table 1 shows, under economic architecture A the industry just barely breaks even at the desired size, while under economic architecture B the industry earns positive profits at the desired size. Figure 1 places the architectures in the program’s sustainable competition diagram, illustrating relevant system cost and government-derived revenue thresholds for different sustainable industry sizes. Note that economic architecture A sits on the boundary between the sustainable duopoly and sustainable monopoly regions, while economic architecture B lies comfortably within the sustainable triopoly region.
@@
Market revenue ($R^{M}_{i}$) \$0.5 B/year/firm \$0.5 B/year/firm
@@
Both representations show that under economic architecture A, high-cost shocks or low-market demand shocks are likely to require additional budget outlays to prevent the industry from collapsing to a monopoly. Equivalently, the industry can weather greater uncertainty under economic architecture B than under economic architecture A. In the next section, I develop an application to illustrate how this framework can be used to support program design.
3 An application to commercial crewed space stations
@@
Commercial space stations exhibit the high fixed costs, uncertain market demand, and opportunities for shared infrastructure that characterize many advanced space technology systems. Suppose a public space agency with a human spaceflight program intended to maintain continuous human presence in low-Earth orbit while developing sustainable commercial markets for space station services.777“Continuous presence” has been a stated goal of U.S. space policy for some time (NASA, 2024). Such non-economic goals are common in space policy, which is often meant to support a sense of national achievement. MacDonald (2017) discusses this dimension of space economics in greater detail. Should the agency support multiple independent “free-flyer” stations, each providing complete space station capabilities, or should it invest in shared infrastructure that multiple firms can utilize while developing their own specialized modules? Using cost estimates derived from analogous system elements and budget assumptions consistent with NASA planning, I compare these two approaches to demonstrate the framework’s application to program design decisions.
@@
A program supporting continuous human presence requires substantial investment in both station capabilities and crew operations, with private investors expecting returns over the asset’s full lifetime. I assume the program covers transportation to and from the station as well as crew supplies and equipment. Since transportation and crew support costs remain roughly constant regardless of whether the program supports independent stations or shared infrastructure, the economically relevant comparison focuses on the “capturable budget”—funds available for direct purchases, direct transfers, and shared infrastructure investment. Based on the 2026 NASA President’s Budget Request for commercial space station services, this capturable budget amounts to approximately \$1B annually. [^8] The 2026 NASA President’s Budget Request contains roughly \$2B for crewed space station services in 2030, when the ISS is retired and private stations become operational (NASA, 2025a). This includes approximately \$1.2B for transportation and resupply flights. The remaining \$800M represents the capturable portion available for station-related investments, rounded to \$1B for analytical convenience. I assume market demand from research institutions, private companies, and other national space agencies and government entities provides an additional \$500M annually in total potential revenue for station operators.
@@
The economic architecture choices regarding shared infrastructure stem from the separability of space station functions. Space stations provide two distinct types of capabilities: core functions that every crewed station requires, and habitat functions that differentiate market offerings. Core functions include life support systems capable of sustaining crew between resupply missions, power generation and distribution, command and data handling, and docking capabilities for crew and cargo vehicles. These systems must meet stringent safety and performance requirements regardless of the station’s intended market—a facility serving government researchers needs the same fundamental life support capabilities as one targeting commercial customers. [^9] The Environmental Control and Life Support System (ECLSS) on the ISS currently performs at about 42% O2 recovery from CO2 air loop closure (NASA, 2022). A station aimed at long-duration research missions may require higher recovery rates than one serving short-duration tourists, but both need reliable life support systems. Habitat functions, by contrast, determine what activities crew can perform: research laboratories, manufacturing facilities, recreational amenities, or specialized equipment deployment capabilities. A crewed station operator targeting pharmaceutical research will invest differently in habitat capabilities than one serving space tourism, but both require identical core infrastructure to keep humans alive and connected to Earth. This functional separation creates a natural opportunity for shared infrastructure investment, where the public agency provides essential core capabilities that all operators need, while private firms differentiate their offerings through specialized habitat modules.
@@
Suppose the program supports continuous presence through stations accommodating up to 8 crew members, with government flights carrying up to 2 agency astronauts per flight through two crew rotations annually, leaving at least 2 seats per flight available for other paying customers. [^10] I am not assuming the government space agency fully fills each flight with its own astronauts, only that the agency pays for the flight. The public space agency may use the seats to barter with other space agencies, to sell back to the public at a discount, or find another use for them. Barter arrangements between space agencies using the ISS are not uncommon (Veldhuyzen and Grifoni, 1999; Selding, 2014; Foust, 2025). This flight cadence, supported by two dedicated resupply missions annually, ensures substantial station capacity remains available for non-government customers, creating the market opportunity that private investors must evaluate. I assume stations have 15-year design lifetimes with construction beginning in 2025, and that private investors expect 5% annual returns over the asset lifetime. [^11] Depreciation and maintenance requirements will depend on the system architecture; I assume these are captured in the operating costs. I assume these costs and revenues are fixed in base year terms for the duration of the program. In the independent free-flyer architecture, each firm develops complete station capabilities including both core and habitat functions. In the shared core module architecture, the government provides a common core infrastructure supporting life support, power, docking, and robotic arm capabilities, while private firms develop specialized habitat modules that connect to this shared foundation. In the shared core architecture, I assume the core module can accommodate up to two habitat modules, with government demand for station services split equally between providers in scenarios with multiple competitors. These assumptions are consistent with current NASA use of the ISS and the capabilities of existing crew and cargo vehicles.

### v2

@@ Abstract
Advanced space technology systems often face high fixed costs, can serve limited external demand, and are significantly driven by non-economic motivations. While increased entrepreneurial activity and national ambitions in space have encouraged planners at public space agencies to develop markets around such systems, the very factors that make the recent growth of the space economy so remarkable also challenge planners’ efforts to develop and sustain markets for space-related goods and services. This paper proposes a diagram to visualize the number of competitors a market can sustain as a function of the industry’s cost structure; the distribution of program support across direct purchases, direct investments, and shared infrastructure; and the magnitude of demand external to the program. Building on public goods theory, the diagram shows how marginal dollars invested in shared infrastructure can create non-rival benefits supporting more competitors than direct purchases or subsidies. The diagram is demonstrated with a stylized application inspired by NASA’s Commercial LEO Destinations program. Under conditions consistent with public data, independent stations generate industry-wide losses of \$355 million annually, while shared core infrastructure enables industry-wide profits of \$154 million annually. The choice between shared infrastructure and direct purchases can depend on economic conditions that are outside the agency’s control. Under favorable conditions—e.g., strong demand and low capital costs—direct purchases may suffice to sustain competition. Under challenging conditions—e.g., limited demand or tight capital markets—even well-designed shared infrastructure may prove insufficient to sustain competition.
@@ 1 Introduction
Public space agencies are increasingly interested in developing sustainable competition to supply advanced space technologies to meet growing commercial and international interests. Several have adopted market development goals, in some cases seeking to expand their commercial procurement of space services while maintaining certain national industrial capabilities. [^1] For example, public space agencies in the U.S., Japan, and Europe have all announced various types of market development goals (NASA, 2019; European Commission, 2025; ESA, 2025a; Cabinet Office, Government of Japan, 2024; EY Global, 2024), while NASA’s investment in the Space Launch System was influenced by industrial base considerations (NASA, 2011). As space ambitions rise globally while public budgets face greater strain, understanding and clearly communicating principles of pro-competitive public investment in advanced technology systems is an increasingly important task for economists and public space agency officials. Yet a basic economic question for market development—how should a program allocate its budget across purchases, transfers, and infrastructure to sustain competition to supply space-related goods and services—remains underexplored.
Space market development efforts face deep uncertainty. Measurements of the current space economy are contested, with independent analyses revealing that some industry projections systematically overstate growth by double-counting government revenues and input costs (Crane et al., 2020). Forecasting demand for markets that don’t yet exist—like lunar surface power or commercial space stations—is more challenging still. Cost structures are often speculative, and beliefs about the relative efficiency of different contract types often lack rigorous econometric grounding. Factors like external demand (i.e., demand outside the agency’s control) and interest rates can overwhelm program design choices, leaving agencies vulnerable to conditions beyond their control.
Despite this uncertainty, public space agencies must decide how to distribute program budgets across direct purchases, transfers, and shared infrastructure. This paper introduces a “sustainable competition” diagram showing how many firms can be sustained at different levels of program spending, external demand, and cost imposed by requirements, to help planners structure thinking about economic trade-offs and uncertainty in program design. The diagram shows boundaries between profitability regions as a function of the agency’s choice variables, making it easy to assess program design considerations like whether particular expenditures can sustain the desired number of firms and how large a shock would be required to threaten the program’s economic sustainability. Building on the standard theory of public goods (Samuelson, 1954), the key insight is that marginal dollars invested in shared infrastructure can create non-rival benefits supporting more competitors than direct purchases or subsidies.
Space situational awareness illustrates this logic. The U.S. Space Force operates the Space Surveillance Network, a system of ground-based sensors and satellites tracking more than 49,000 objects in Earth orbit (18th Space Defense Squadron, 2024). The 18th Space Defense Squadron maintains this catalog and makes it freely available to U.S. and international satellite operators, academia, and other entities through space-track.org in order to reduce the risk of collisions between satellites. Public provision of basic tracking services as shared infrastructure lowers the cost of entry for commercial satellite operators, facilitating market development. Private firms can use the catalog to provide value-added analytics services without recreating the full network of ground-based sensors and satellites.
@@
A stylized application based on NASA’s Commercial LEO Destinations program formalizes this logic and illustrates use of the diagram. The application compares independent free-flyer versus shared core module architectures for sustaining at least two competing crewed space station operators. Under cost and demand conditions consistent with public data, independent stations generate industry-wide losses of \$355 million annually, while shared core infrastructure enables industry-wide profits of \$154 million annually.
Section 2 describes the general approach assessing economics of market development programs and constructing the diagram. Section 3 develops estimates and constructs the diagram for an application to crewed space stations in low-Earth orbit. Section 4 identifies key directions for future research on public investment and market development strategies for advanced technologies. Section 5 concludes.
2 A diagram for assessing economics of market development programs
@@
Consider a public space agency with annual budget $B$ designing a program to develop sustainable markets for a class of advanced space technology systems. These systems—such as orbital debris tracking radars, lunar surface power sources, or orbital debris removal spacecraft—typically exhibit high fixed costs, long development timelines, and highly uncertain external demand, creating conditions where uncoordinated private investment may result in too few competitors or none at all. As Weinzierl (2018) notes, complementarities between space technologies mean space businesses across the technology stack play stag hunts with each other. This analysis assumes the public space agency is a planner capable of using its program requirements and budget to structure technological complementarities between private actors to encourage desired outcomes. The $N$ firms that develop and use the technology systems of interest are referred to as “the industry.” All economic variables are expressed as equivalent annuity units to facilitate comparison across different payment timing structures, firms are assumed to be symmetric in costs and capabilities, and inflation is ignored.
@@
Public investment into shared infrastructure provides benefits $Y^{G}$ to the industry according to the valuation function $f(G^{S})$, which measures how much shared infrastructure reduces firms’ costs relative to developing equivalent capabilities independently. [^3] The specific functional form depends on the technology systems and business models involved. For modular systems, benefits may be approximately linear up to shareable essential element costs, then exhibit diminishing returns; for network systems requiring interoperability, benefits may show increasing returns to scale. Here, firms are assumed to uniformly value shared infrastructure at cost: $Y^{G}=G^{S}$.
Firms in the industry earn total revenue $R_{i}$, with $R_{i}^{G}$ earned from the program and $R^{M}_{i}$ earned from sources external to the focal agency—individuals, households, firms, other public space agencies, and any other entities that may purchase the industry’s output. This separates revenues the focal agency controls from those it cannot directly control; while the agency may coordinate with other entities, the agency does not control their spending. Under symmetry, each firm receives equal shares: $R^{G}_{i}=R^{G}/N$ and $R^{M}_{i}=R^{M}/N$. Similarly, direct transfers simplify to $G_{i}=G^{D}/N$.
Total industry revenue is $R=\sum_{i}(R_{i}^{G}+R_{i}^{M})$. Firms incur gross total cost $X_{i}$ and net total cost $C_{i}=X_{i}-Y^{G}-G_{i}$; industry total cost is $C=\sum_{i}C_{i}$. The gross costs $X_{i}$ represent costs to develop, deploy, and operate the target system; under symmetry, $X_{i}=X/N$ where $X=\sum_{i}X_{i}$ is industry-level gross total cost. [^4] Space technology program managers recognize a positive relationship between system performance and cost (NASA Executive Cost Analysis Steering Group, 2015), allowing cost to proxy—with suitable caveats—for system performance required by the program.
@@
A marginal dollar provides greater support to the industry if invested in shared infrastructure than if used for direct purchasing when:
@@
If the public space agency wants to maintain at least $N$ competitors, then it must ensure $R^{G}$, $G^{D}$, and $Y^{G}$—its support for the industry—are sufficient to ensure all $N$ firms earn non-negative profits, conditional on the anticipated level of external demand, $R^{M}$. Formally:
@@
This constraint captures the fundamental trade-off: supporting more competitors requires either reducing individual firm costs (through shared infrastructure) or increasing industry revenues (through direct purchases, transfers, or market growth). Programs violating this constraint will see firms exit, potentially ending in monopoly or no suppliers. [^5] Questions regarding payment timing and conditioning can be investigated by appropriately modeling and discounting benefit flows when calculating $R^{G}$, $G^{D}$, and $Y^{G}$.
@@
Overlaying profitability regions across different industry sizes shows the full architectural possibility set. For fixed $N$, these regions are $\{(R^{G},C):~\forall i,~\Pi_{i}(N)\geq 0\}$—intersections of superlevel sets for firms’ profitability constraints. [^6] Similar diagrams are used in managing engineering costs and schedules in advanced technology programs, e.g., Joint Confidence Level plots (NASA Executive Cost Analysis Steering Group, 2015). Direct government purchases, $R^{G}$, represent agency support appearing on firms’ income statements; industry total costs, $C=X-Y^{G}-G^{D}$, show the cost burden firms face after accounting for shared infrastructure benefits and direct transfers. Regions where profits are non-negative show combinations of costs and purchases that can sustain a given number of firms.
To illustrate, consider two architectures for sustaining at least two firms: architecture A transfers all funds as non-dilutive grants; architecture B divides its budget evenly between direct purchases and shared infrastructure. With external revenues of \$1B/year, gross costs of \$2B/year, program budget of \$1B/year, and firms valuing shared infrastructure at cost. Table 1 lists the economic parameters. Under architecture A the industry barely breaks even; under architecture B it earns positive profits. Figure 1 shows architecture A on the boundary between sustainable duopoly and monopoly regions, while architecture B lies within the sustainable triopoly region.
@@
External revenue ($R^{M}_{i}$) \$0.5 B/year/firm \$0.5 B/year/firm
@@
Both representations show that architecture A requires additional budget outlays following cost or demand shocks to prevent industry collapse to monopoly, while architecture B can weather greater uncertainty. The next section illustrates how to construct the diagram with an application.
3 An application to crewed space stations
@@
Crewed space stations exhibit the high fixed costs, uncertain external demand, and opportunities for shared infrastructure that characterize many advanced space technology systems. Like the tracking catalog in the Space Situational Awareness case, crewed stations also have foundational infrastructure that can be shared across firms. Suppose a public space agency with a human spaceflight program intended to maintain continuous human presence in low-Earth orbit while developing sustainable markets for crewed space station services.777“Continuous presence” has been a stated goal of U.S. space policy for some time (NASA, 2024b). While a classical economic planner may focus on goals derived from economic welfare maximization, non-economic goals are common in space policy, which is often meant to support a sense of national achievement. MacDonald (2017) discusses this dimension of space economics in greater detail. Should the agency support multiple independent “free-flyer” stations, each providing complete space station capabilities, or should it invest in shared infrastructure that multiple firms can use while developing their own specialized modules? Using cost estimates derived from analogous system elements and budget assumptions consistent with NASA planning, this application compares these two approaches to demonstrate the diagram’s use in program design decisions.
@@
A program supporting continuous human presence requires substantial investment in both station capabilities and crew operations, with private investors expecting returns over the asset’s full lifetime. The analysis assumes the program covers transportation to and from the station as well as crew supplies and equipment. Since transportation and crew support costs remain roughly constant regardless of whether the program supports independent stations or shared infrastructure, the economically relevant comparison focuses on the “capturable budget”—funds available for direct purchases, direct transfers, and shared infrastructure investment. Based on the 2026 NASA President’s Budget Request for commercial space station services, this capturable budget amounts to approximately \$1B annually. [^8] The 2026 NASA President’s Budget Request contains roughly \$2B for crewed space station services in 2030, when the ISS is retired and private stations become operational (NASA, 2025a). This includes approximately \$1.2B for transportation and resupply flights. The remaining \$800M represents the capturable portion available for station-related investments, rounded to \$1B for analytical convenience. The analysis assumes external demand from research institutions, private companies, and other national space agencies and government entities provides an additional \$500M annually in total potential revenue for station operators.
@@
The economic architecture choices regarding shared infrastructure stem from the separability of space station functions. Space stations provide two distinct types of capabilities: core functions that every crewed station requires, and habitat functions that differentiate market offerings. Core functions include life support systems capable of sustaining crew between resupply missions, power generation and distribution, command and data handling, and docking capabilities for crew and cargo vehicles. These systems must meet stringent safety and performance requirements regardless of the station’s intended market—a facility serving government researchers needs the same fundamental life support capabilities as one targeting commercial customers. [^9] The Environmental Control and Life Support System (ECLSS) on the ISS currently performs at about 42% O2 recovery from CO2 air loop closure (NASA, 2022). A station aimed at long-duration research missions may require higher recovery rates than one serving short-duration tourists, but both need reliable life support systems. Habitat functions, by contrast, determine what activities crew can perform: research laboratories, manufacturing facilities, recreational amenities, or other specialized capabilities. A crewed station operator targeting pharmaceutical research will invest differently in habitat capabilities than one serving space tourism, but both require similar core infrastructure to keep humans alive and connected to Earth. The core functions represent foundational infrastructure—essential prerequisites without which crewed operations cannot occur—while habitat capabilities represent enabling infrastructure that expands opportunities for service differentiation. This functional separation creates a natural opportunity for shared infrastructure investment, where the public agency provides essential core capabilities that all operators need, while private firms differentiate their offerings through specialized habitat modules. This parallels the Space Situational Awareness case: the tracking catalog is foundational infrastructure, and value-added analytics are how Space Situational Awareness providers differentiate their offerings.
@@
Suppose the program supports continuous presence through stations accommodating up to 8 crew members, with government flights carrying up to 2 agency astronauts per flight through two crew rotations annually, leaving at least 2 seats per flight available for other paying customers. [^10] The analysis does not assume the government space agency fully fills each flight with its own astronauts, only that the agency pays for the flight. The public space agency may use the seats to barter with other space agencies, to sell back to the public at a discount, or find another use for them. Barter arrangements between space agencies using the ISS are not uncommon (Veldhuyzen and Grifoni, 1999; Selding, 2014; Foust, 2025). This flight cadence, supported by two dedicated resupply missions annually, ensures substantial station capacity remains available for non-government customers, creating the market opportunity that private investors must evaluate. Stations are assumed to have 15-year design lifetimes with construction beginning in 2025, and private investors are assumed to expect 5% annual returns over the asset lifetime. [^11] Depreciation and maintenance requirements will depend on the system architecture; the analysis assumes these are captured in the operating costs. These costs and revenues are assumed to be fixed in base year terms for the duration of the program. In the independent free-flyer architecture, each firm develops complete station capabilities including both core and habitat functions. In the shared core module architecture, the government provides a common core infrastructure supporting life support, power, docking, and robotic arm capabilities, while private firms develop specialized habitat modules that connect to this shared foundation. The analysis assumes the shared core module can accommodate up to two habitat modules, with government demand for station services split equally between providers in scenarios with multiple competitors. These assumptions are consistent with current NASA use of the ISS and the capabilities of existing crew and cargo vehicles.